OpenStudy (anonymous):

I'm having trouble telling the difference between solution sets. |-3x-12|>-24 Any clue?

OpenStudy (amistre64):

when is an absolute value greater than a negative number?

OpenStudy (amistre64):

is that a -24? or is that spose to be >= 24?

OpenStudy (anonymous):

When it's positive. I understand somewhat what it is supposed to be but my professor is a bit confusing. An problem like.: |-5x-10|>-30 That's \[(-\infty, \infty)\] I need an explanation of why is this so.

OpenStudy (anonymous):

Well rather, why infinity rather than numbers?

OpenStudy (jamesj):

What's the solution to |x| = 0 ? It's x = 0. Now for all x \[ |x| \geq 0 \] Hence if you have to solve \[ |x| \geq 0 \] that's all real numbers. If you have \[ |x-1| \geq 0 \] that's satisfied by all real numbers If you have for some arbitrary constant a \[ |x-a| \geq 0 \] that's satisfied by all real numbers. Hence \[ |x-a| \geq -1 \] is also satisfied by all real numbers, as is \[ |bx-a| \geq -1 \] Now can you see why it's true for your inequality?

OpenStudy (anonymous):

Well let me give you another one and maybe you might be able to explain a little more. \[|2x-6|\ge0\] The solution is \[(-\infty, \infty)\] As opposed to: \[|3x-18|\ge0\] Where \[(-\infty, 6)u(6,\infty)\] Why is this?

OpenStudy (jamesj):

The solution to the second inequality is wrong. \[ |3x - 18| \geq 0 \] if and only if (iff) \[ 3|x-6| \geq 0 \] iff \[ |x-6| \geq 0 \] iff \( x \) is any real number.

OpenStudy (anonymous):

This isn't my answer, this is my professors. I believe I am thinking the same way that you are but this is his way of doing it. He has taught for twenty some years so, I might have to sit down with him and have a conversation on this.

OpenStudy (jamesj):

Ah, I know what the inequality is. You mean this: \[ |3x - 18| > 0 \] strictly greater than zero. Then 6 is not in the solution set.

OpenStudy (anonymous):

Mm no its still |3x−18|≥0

OpenStudy (anonymous):

It is what he wrote on the blackboard.

OpenStudy (jamesj):

Well, either he meant to write > or he's made an error in the solution.

OpenStudy (anonymous):

I'm not entirely sure. Good part about this work is that it isn't graded at all so it wont affect my grade if I don't do them now. I will at least have the opportunity to sit down with him for a moment. I might try to get with him this Sunday afternoon because I know for a fact, Math isn't necessarily my strongest attribute, more of a literature/science guy. My test isn't for another week or so I have sometime. But thank you for helping me! I really appreciate you trying.